Use Intercepts graph equation
2x - 6 = 3y
Are the intercepts (0, -2) Is this the only one.
Can I plot (6,2) as another point?

Answer by rfer(16322)

put into slope intercept form first.
y=mx+b
3y=2x-6
y=2/3x-2
y int b=-2 (0,-2)
slope m=2/3
-------------
y=mx+b
0=2/3x-2
2/3x=2
x=3*3/2
x=6/2
x=3
x int =(3,0)